OpenStudy (anonymous):

a tangent drawn to y=-(x^2 -4) at the point (1, 3) forms a right triangle with the coordinate axes. find the area of this right triangle. please help and show all steps!

OpenStudy (turingtest):

find f'(1) to find the slope of the function at (1,3) then us that value in point-slope form\[y-y_0=f'(x_0)(x-x_0)\]

OpenStudy (turingtest):

that will allow you to fins the intercepts of the line with the coordinate axis, which gives you the length of the sides of the triangle.

OpenStudy (anonymous):

okay so is the slope -2?

OpenStudy (turingtest):

yes f'(1)=-2

OpenStudy (anonymous):

okay so then is the area 16?

OpenStudy (turingtest):

I'm not sure, I haven't tried it, give me a minute...

OpenStudy (turingtest):

I got 25/4 what did you get for the intercepts?

OpenStudy (turingtest):

I got y-3=-2(x-1) y=-2x+5 x-int=5/2 y-int=5 area=[(5/2)*5]/2=25/4

OpenStudy (anonymous):

okay that makes sense thank you!